OpenStudy (anonymous):

Simplify -4|-5| - [2(-17)]

OpenStudy (fibonaccichick666):

@perl

OpenStudy (anonymous):

I know the first part would be -20

OpenStudy (anonymous):

54

OpenStudy (fibonaccichick666):

yes, the first part is -20, please continue @smiley2015

OpenStudy (anonymous):

NO! YOU CAN'T MAKE ME!

OpenStudy (fibonaccichick666):

please do not just give answers @Brostep0s it is against the code of conduct

OpenStudy (anonymous):

NOOOOOOOOOOOOOO!

OpenStudy (anonymous):

14

OpenStudy (anonymous):

−4(|−5|)−(2)(−17) =(−4)(5)−(2)(−17) =−20−(2)(−17) =−20−(−34) =14

OpenStudy (fibonaccichick666):

please do not just give answers @skate-4-life270 it is against the code of conduct